# OpenSlot: Mixed Open-set Recognition with Object-centric Learning

### v1

@@ Abstract
Existing open-set recognition (OSR) studies typically assume that each image contains only one class label, and the unknown test set (negative) has a disjoint label space from the known test set (positive), a scenario termed full-label shift. This paper introduces the mixed OSR problem, where test images contain multiple class semantics, with known and unknown classes co-occurring in negatives, leading to a more challenging super-label shift. Addressing the mixed OSR requires classification models to accurately distinguish different class semantics within images and measure their “knowness”. In this study, we propose the OpenSlot framework, built upon object-centric learning. OpenSlot utilizes slot features to represent diverse class semantics and produce class predictions. Through our proposed anti-noise-slot (ANS) technique, we mitigate the impact of noise (invalid and background) slots during classification training, effectively addressing the semantic misalignment between class predictions and the ground truth. We conduct extensive experiments with OpenSlot on mixed & conventional OSR benchmarks. Without elaborate designs, OpenSlot not only exceeds existing OSR studies in detecting super-label shifts across single & multi-label mixed OSR tasks but also achieves state-of-the-art performance on conventional benchmarks. Remarkably, our method can localize class objects without using bounding boxes during training. The competitive performance in open-set object detection demonstrates OpenSlot’s ability to explicitly explain label shifts and benefits in computational efficiency and generalization.
@@ I Introduction
Nevertheless, real images are often multi-labeled, with mixed known and unknown class objects. In this study, we formulate a mixed OSR problem (Fig. 1, (b)) that considers test samples with multiple class semantics, which is more commonly seen in real life. In each negative image, the known and unknown classes occur jointly. In this setting, the class set of negative images is the super-set of positives. Therefore, we define the mixed OSR as the super-label shift.
@@
Motivated by these advances [33, 35], we present a novel OpenSlot framework, propose learning slot features to represent the different class semantics, and implement image classification to handle the mixed OSR. To isolate the invalid and suspicious background slots during classification, we specifically propose a simple yet effective anti-noise-slot (ANS) technique to address the semantic misalignment problem of the class prediction. Combined with out-of-distribution (OOD) detection scoring metrics, we measure the “knownness” of the object semantic that each slot corresponds to, and achieve significant improvements on various OSR benchmarks. Additionally, our method could explicitly detect unknown objects without using bounding boxes for training, achieving competitive results in open-set object detection.
@@
  • We introduce the mixed OSR task, which requires detecting unknown class objects in images with known content.
@@ III Mixed Open-set recognition
In this setting, we use $y_{i}$ to denote $\mathbf{x_{i}}$’s class label set and get $y_{i}\cap\mathcal{K}\neq\emptyset\quad\text{and}\quad y_{i}\cap\mathcal{A}\neq\emptyset$. Following this setup, we get $\mathcal{Y}=\mathcal{A}\cup\mathcal{K}$ and see that the distribution shift happens in the superset of $\mathcal{K}$ (termed super-label shift).
Handling this mixed problem necessitates OSR models to recognize all class objects within the negative sample $\mathbf{x_{i}}$ and group them as known or unknown. Besides, detecting the super-label shift in images with multiple class semantics is crucial for extending OSR to multi-label scenarios; a larger label space implies more mixture scenarios, with distribution shifts potentially manifesting anywhere. This capability directly benefits real-world applications, such as autonomous driving, species discovery, medical imaging analysis [14], and defect detection [15] in manufacturing processes.
Intuitively, classifying an image with multiple semantics can be significantly influenced by the known class objects. One potential solution is to use pre-trained dense predictors, e.g., object detecter [32] and segmentation methods [19], to exhaustively classify each object proposal and pixel for precise identification. However, this approach needs more label information in training and can be computationally expensive (compared in Sec. V-E). Besides, their OSR performance relies heavily on dense predictors, which suffer from domain shifts [30]. These challenges motivate us to develop a classification model to obtain generalizable representations of diverse semantics within images and learn clear decision boundaries among different category groups to perceive label shifts.
@@ IV Method
Figure 2: OpenSlot overview. In (1), we pre-train Dinosaur [33] (following the encoder-decoder design, the encoder can be either ResNet or Transformer architecture) on the known training images, using slots $S$ to learn class semantics. Next (2), we add two classifiers, the noisze classifier $\varphi_{nz}$ and the main classifier $\varphi_{fg}$, trained with $L_{nz}$ (Eq. 2) and $L_{hun}$ (Eq. 4) respectively, to identify the noise (invalid and background) slots and implement semantic classification with the proposed ANS technique (Fig. 3). During inference (3), we score slot-based class predictions to address the mixed OSR.
@@ IV-A Preliminary: Object-centric learning
The goal of OCL [28, 18] is to represent the visual input with a set of object features (termed slots), with which to obtain the compositional representation of the input to facilitate downstream vision tasks [43, 17]. In training, slot are randomly initialized and refined by an iterative attention module [28], to capture an independent object semantic.
OCL commonly adopts the autoencoder design and uses input reconstruction as the objective function. Recently, Dinosaur [33] proposes reconstructing the self-supervised features, e.g., DINO [1], during training and scales OCL to natural images, achieving the first success in the unsupervised setting. When combining the Transformer decoder [35], Dinosaur represents objects belonging to the same class with one slot, showing the potential of slot-based image classification.
Motivations. We propose utilizing slot features to perform multi-class semantic classification and address the mixed OSR.
Firstly, each slot represents a particular object semantic and produces an independent class prediction. When evaluating the “knowness” of a negative image, the slot representing unknown classes (UUCs) contributes a significant score value difference compared to known classes. In contrast, existing methods [38, 37] typically aggregate all semantics to obtain an overall class prediction. Consequently, the resulting OSR score can be largely influenced by the known classes (KKCs) present in the input (Section V-D compares our slot-based method with existing studies regarding OSR scoring).
Secondly, slots inherently carry spatial location information [28, 33] of the represented objects, enabling our work to localize unknown object semantics and display the super-label shift of the test sample explicitly. Remarkably, relying solely on class labels during training, our method is capable of performing open-set object detection (OSOD) and achieves competitive performances (see Sec. V-E).
@@ IV-B OpenSlot
For each $I\in\mathcal{D}_{train}$ ($I\in\mathbb{R}^{H\times W\times 3}$), we get a set of slots $S=\{S_{1},...S_{N}\}$ and their attention masks $A$ (Fig. 2 (1)):
@@
However, the number of slots $N$ is fixed during training; OCL learns invalid slots [28] when $N$ is larger than that of practical semantics. Meanwhile, the slots that represent background information also increase the difficulty of matching. These two issues would cause a severe misalignment problem, i.e., the final classification decision does not actually represent the class object content of the image. To address these limitations (shown and ablated in Sec. V-C), we propose the anti-noise-slot (ANS) technique to mitigate the effects of noise (invalid and background) slots on image classification.
Firstly, we add an MLP classifier $\varphi_{fg}$ (shared across slots) after the slot-attention encoder to implement class prediction and matching. To isolate the invalid slots, we consider the attention masks $A$ (shown in Fig. 3) a vital cue that provides rich information for the separation. We use a threshold $\alpha$ to determine the semantic region of $A$, and assume that to denote a valid object semantic (foreground/background unknown), the slot’s attention mask should have over two consecutive pixels (corresponding to patch features) over $\alpha$. In this step, we use a mask $M_{inv}\in\mathbb{R}^{N\times 1}$ for the invalid slot separation.
@@
Let $L_{set}=CE(\varphi_{fg}(S),\mathbf{L})$ be the pairwise loss between the slot predictions $\varphi_{fg}(S)$ and the label set $\mathbf{L}$, where CE is cross-entropy loss. With the noise mask information from $M$, we recalibrate the assignment cost of Hungarian matching as:
$L_{hun}=(1-M)\cdot L_{set}+\lambda M\cdot L_{set},$ (4)
@@
Figure 3: Anti-noise-slot (ANS). After pre-training, we learned slot $S$ to (a) represent the input’s diverse semantics (shown in different colors) and get their attention masks (each slot has an attention mask). (b) In each step of the classification training, we: 1) first merge the null slot mask with the invalid slot mask $M_{inv}$ to get the pseudo label $M_{nz}$; (2) use $M_{nz}$ to train the noise classifier $\varphi_{nz}$; 3) get the high-confidence prediction $M$ from $\varphi_{nz}$ to calibrate (Eq. 4) the assignment cost of Hungarian matching, enabling $\varphi_{fg}$ to focus on true class semantics.
[…]
Framework Overview. As illustrated in Fig. 2, we firstly pre-train Dinosaur [33] on $\mathcal{D}_{train}$ (the training set of known classes) to obtain the representative slot features. Next, we frozen the Dinosaur part and added two classifiers, $\varphi_{nz}$ and $\varphi_{fg}$, following the encoder. After excluding the noise slots from the semantic matching, we implement the image classification. The overall loss is expressed as follows:
$L=L_{hun}+L_{nz}.$ (5)
@@
Inference. To get the classification result in the closed set, we use the max-logit slot in single-label experiments and the maximum class-wise logit across all slots in multi-label cases.
For OSR tasks, we score slot predictions with common OOD detection functions [27, 23, 13, 24]. The score measures the label shift degree of the content that each slot corresponds to. Lastly, we sum all scores as the decision score.
@@ V-B Results of OSR
Single-label. Tab. I compares OpenSlot with existing OSR methods, including CSSR [16], ARPL [2], MLS [37], and a transformer baseline VIT-MLS. First, all baselines experience a substantial accuracy degradation on the mixed test set $\mathcal{M}$ (compared to $\mathcal{H}$). This indicates that the super-label shift in the mixed OSR is more challenging than the conventional full-label shift, revealing the existing models’ limitations in processing mixed semantics.
@@ V-C Ablation studies
Effects of ANS. By comparison (Fig. 4, ours), ANS ensures that the max-logit slot falls into the FG group, aligning with the true class semantics. Furthermore, ANS increases the logit value gap between the FG and Noise slot groups.
Meanwhile, we perform a similar sorting and visualization analysis of the attention masks in Fig. 6. On the right side, the main classifier $\varphi_{fg}$ suppresses noise and assigns the slot representing the true class object with the largest logit value. This finding is consistent with Fig. 4 (ours). On the left figure, the sorted slots based on the outputs from the noise classifier $\varphi_{nz}$ showcase that $\varphi_{nz}$ accurately discriminates noise features.
Fig. 4 and the comparison between Fig. 5 and 6 showcase that our ANS method effectively separates the noise slots and excludes them from the class classification. During inference, when computing OSR scores for positive and negative images, the noise slots of both images would output similar values. This ensures that the primary score value gap between the positive and the negative stems from the semantic difference of class objects, leading to a clear class decision boundary.
Finally, we quantitatively compare Pure Slot and our OpenSlot in Tab. IV. Our consistent performance gains achieved on both the $\mathcal{M}$ and $\mathcal{H}$ test sets across two benchmarks verify the benefits of ANS in enhancing the capability of slots to detect both super-label and full-label shifts.
Choice of scoring functions. To obtain the OSR results, we score the slot-based class predictions. We evaluate various out-of-distribution (OOD) detection methods, including four logit-based: MSP [13], ODIN [24], Energy [27], MaxLogit [37], as well as the feature-based Mahalanobis distance [23]. In Tab. V, we compare their performances. When combined with OpenSlot, MaxLogit and Energy [27] consistently outperform other scoring methods, yielding superior OSR results.
@@
m0.3cm¡m0.4cm¡—m 1.4cm¡m 1.4cm¡m 1.4cm¡m 1.4cm¡ $\alpha$ $\beta$ VOC-6/14 COCO-20/60
@@ V-D Effect of slot prediction in OSR scoring
By contrast, our slot approach addresses the threat of known classes in the mixed OSR problem by producing independent class predictions for captured image semantics. During inference, every slot prediction in $f_{slot}$ is measured with the score function and is counted for OSR evaluation, without effects from others. In (Fig. 7, bottom right), we observe that the score gap between the positive and the negative image primarily comes from the detected unknown class object.
@@
V-E Yielding Benefits to Open-Set Object Detection
Results. Table VIII compares OpenSlot with existing OSOD studies. Without employing a Region Proposal Network [32], OpenSlot even outperforms several object detector-based approaches and achieves performance comparable to SIREN-vMF [7]. Fig. 8 presents example results where our method corrects the mispredictions of the object detector, reducing the false positive rate.
Notably, unlike traditional methods that rely on predefined classes and selective search algorithms for proposal generation, OpenSlot discovers objects from image data in an unsupervised manner [1, 33], and can obtain robust object representations against domain shifts [5]. When tested on open-set images with significant domain differences from the closet-set (Fig. 9), these properties ensure that OpenSlot identifies different objects of interest, leading to improved localization of unknown objects. and reflecting our slot approach’s stronger generalization ability than detector-based methods.
Overhead comparison. Tab. IX compares the computational cost of OpenSlot and Faster R-CNN [32] from different perspectives. We have a lightweight and with only 2.4 million parameters (including the pre-train and the classification models), and require merely 3.5 hours for training. This cost is significantly lower than object detector models. Moreover, OpenSlot demonstrates faster inference time when generating object detection results, highlighting its efficiency advantages.
@@ V-E Yielding Benefits to Open-Set Object Detection
Method Param. Train. time Infer. time
Faster-RCNN [32] 33.0 M 3.7 h 0.26 s
Ours 2.4 M 3.5 h 0.21 s
TABLE IX: We compare Faster-RCNN and OpenSlot on RTX3090, in terms of the parameter size (million, M), training time (hours, h), and inference speed (second/per image).

### v2

@@ Abstract
Existing open-set recognition (OSR) studies typically assume that each image contains only one class label, with the unknown test set (negative) having a disjoint label space from the known test set (positive), a scenario referred to as full-label shift. This paper introduces the mixed OSR problem, where test images contain multiple class semantics, with both known and unknown classes co-occurring in the negatives, leading to a more complex super-label shift that better reflects real-world scenarios. To tackle this challenge, we propose the OpenSlot framework, based on object-centric learning, which uses slot features to represent diverse class semantics and generate class predictions. The proposed anti-noise slot (ANS) technique helps mitigate the impact of noise (invalid or background) slots during classification training, addressing the semantic misalignment between class predictions and ground truth. We evaluate OpenSlot on both mixed and conventional OSR benchmarks. Without elaborate designs, our method not only excels existing approaches in detecting super-label shifts across OSR tasks, but also achieves state-of-the-art performance on conventional benchmarks. Meanwhile, OpenSlot can localize class objects without using bounding boxes during training, demonstrating competitive performance in open-set object detection and potential for generalization.
@@ I Introduction
Nevertheless, real images are often multi-labeled, with mixed known and unknown class objects. In this study, we formulate a mixed OSR problem (Fig. 1, (b)) that considers test samples with multiple class semantics, which is more commonly seen in real life. In each negative image, the known and unknown classes occur jointly. In this setting, the class set of negative images is the super-set of positives. Therefore, we define mixed OSR as a super-label shift.
@@
Motivated by these advances [36, 39], we present a novel OpenSlot framework for learning slot features to represent different class semantics and implementing image classification to handle the mixed OSR. To isolate the invalid and suspicious background slots during classification, we propose a simple yet effective anti-noise-slot (ANS) technique to address the semantic misalignment problem of the class prediction. Combined with out-of-distribution (OOD) detection scoring metrics, we measure the “knownness” of the object semantic that each slot corresponds to, and achieve significant improvements on various OSR benchmarks. Additionally, our method could explicitly localize unknown objects without using bounding boxes for training, achieving competitive results in open-set object detection tasks.
@@
  • We introduce the mixed OSR, a classification task that requires models to detect unknown class objects in images containing known content.
@@ III Mixed Open-set recognition
In this setting, we use $y_{i}$ to denote $\mathbf{x_{i}}$’s class label set, and have $y_{i}\cap\mathcal{K}\neq\emptyset\quad\text{and}\quad y_{i}\cap\mathcal{A}\neq\emptyset$. Hence, the label space becomes $\mathcal{Y}=\mathcal{A}\cup\mathcal{K}$, and the distribution shift occurs in the superset of $\mathcal{K}$, termed super-label shift.
Handling this mixed problem requires OSR models to recognize all class objects within the negative sample $\mathbf{x_{i}}$ and classify them as known or unknown. Furthermore, detecting the super-label shift in images with multiple class semantics is crucial for extending OSR to multi-label scenarios; a larger label space implies more mixed scenarios, with distribution shifts potentially manifesting anywhere. This capability is directly beneficial for real-world applications, such as autonomous driving, species discovery, medical imaging analysis [16], and defect detection [17] in manufacturing processes.
Intuitively, classifying an image with multiple semantics can be influenced by the presence of known class objects. One potential solution is to leverage pre-trained dense predictors, such as object detectors [34] and segmentation methods [21], to implement open-set object detection [43, 35] and open-set image segmentation [13], classifying each object proposal and pixel for more accurate known/unknown identification. However, these dense approaches require extensive labeled data, are computationally expensive (as discussed in Sec. V-E), and are sensitive to domain shifts [32].
In contrast, our mixed OSR approach relies solely on class labels during training. It perceives label shifts in multi-label and mixed scenarios by learning generalizable representations of diverse semantics within images and establishing clear decision boundaries between known and unknown categories.
@@ IV Method
Figure 2: OpenSlot overview. In (1), we pre-train Dinosaur [36] (following the encoder-decoder design, the encoder can be either ResNet or Transformer architecture) on the known training images, using slots $S$ to learn class semantics. Next (2), we add two classifiers, the noise classifier $\varphi_{nz}$ and the main classifier $\varphi_{fg}$, trained with $L_{nz}$ (Eq. 3) and $L_{match}$ (Eq. 5) respectively, to identify the noise (invalid and background) slots and implement semantic classification with the proposed ANS technique (Fig. 3). During inference (3), we score slot-based class predictions to address the mixed OSR.
@@ IV-A Preliminary: Object-centric learning
The goal of OCL [30, 20] is to represent the visual input with a set of object features (termed slots), with which to obtain the compositional representation of the input to facilitate downstream vision tasks [47, 19]. In training, slots are randomly initialized and refined by an iterative attention module [30], to capture an independent object entity through semantic grouping [10].
OCL commonly adopts the autoencoder design and uses input reconstruction as the objective function. Recently, Dinosaur [36] proposes reconstructing the self-supervised features, e.g., DINO [2], during training and scales OCL to natural images, achieving the first success in the unsupervised setting. This study also shows that the Transformer decoder [39, 38] groups objects of the same class into a single slot, aligning with class-level semantics. Our work chooses Dinosaur because it requires no extra supervision [10] and effectively learns slot-based OCL representations for class semantics, making it well-suited for OCL pretraining and image semantic classification, while leaving other alternative models [37] for future exploration.
Motivations. We propose using slot features to capture diverse image semantics and implement class classification for known/unknown discrimination, thus addressing the mixed OSR. Our approach is motivated by two key advantages:
Firstly, each slot represents a distinct object semantic and produces an independent class prediction. When evaluating the “knowness” of a negative image, the slot representing unknown classes (UUCs) contributes a significant score value difference compared to known classes. In contrast, existing methods [42, 41] typically aggregate all semantics to obtain an overall class prediction. Consequently, the resulting OSR score can be largely influenced by the known classes (KKCs) present in the input (Section V-D compares our slot-based method with existing studies regarding OSR scoring).
Secondly, slots inherently encode spatial location information [30, 36] of the represented objects, enabling our method to localize unknown object semantics and display the super-label shift of the test sample explicitly. Notably, relying solely on class labels during training, our method can perform open-set object detection (OSOD) and achieves competitive performances (see Sec. V-E).
@@ IV-B OpenSlot
For each training image $I\in\mathcal{D}_{train}$ ($I\in\mathbb{R}^{H\times W\times 3}$), we get a set of slots $S=\{S_{1},...S_{N}\}$ and their attention masks $A$ (Fig. 2 (1)):
@@
However, the number of slots $N$ is fixed during training; OCL learns invalid slots [30] when $N$ is larger than that of practical semantics. Meanwhile, the slots that represent background information also increase the difficulty of matching. These two issues would cause a severe misalignment problem, that is, the final classification decision does not actually represent the class object content of the image. To address these limitations (shown and ablated in Sec. V-C), we propose the anti-noise-slot (ANS) technique to mitigate the effects of noise (invalid and background) slots on image classification.
Firstly, we add an MLP classifier $\varphi_{fg}$ (shared across slots) after the slot-attention encoder to implement class prediction and matching. To isolate invalid slots, we consider the attention masks $A$ (shown in Fig. 3) a vital cue that provides rich information for separation. We use a threshold $\alpha$ to determine the semantic region of $A$, and assume that to denote a valid object semantic (foreground/background agnostic), the slot’s attention mask should have over two consecutive pixels (corresponding to patch features) over $\alpha$. In this step, we use a mask $M_{inv}\in\mathbb{R}^{N\times 1}$ for the separation of the invalid slots:
$M_{inv}=\mathbf{1}[A>\alpha],$ (2)
@@
Let $L_{set}=CE(\varphi_{fg}(S),\mathbf{L})$ be the pairwise loss between the slot predictions $\varphi_{fg}(S)$ and the label set $\mathbf{L}$, where CE is cross-entropy loss. With the noise mask information from $M$, We first recalibrate the assignment cost, and then apply the Hungarian algorithm to find the matching with the least cost:
$L_{match}=HUN((1-M)\cdot L_{set}+\lambda M\cdot L_{set}),$ (5)
@@
Figure 3: Anti-noise-slot (ANS). After pre-training, we learned slot $S$ to (a) represent the input’s diverse semantics (shown in different colors) and get their attention masks (each slot has an attention mask). (b) In each step of the classification training, we: 1) first merge the null slot mask with the invalid slot mask $M_{inv}$ (Eq. 2) to get the pseudo label $M_{nz}$; (2) use $M_{nz}$ to train the noise classifier $\varphi_{nz}$; 3) get the mask $M$ (Eq. 4) of high-confidence noise prediction from $\varphi_{nz}$ to calibrate (Eq. 5) the assignment cost of Hungarian matching, enabling $\varphi_{fg}$ to focus on true class semantics.
As shown in Fig. 3, masks $M$ and $M_{nz}$ are constructed based on the prediction of $\varphi_{nz}$ and $\varphi_{fg}$, respectively. These masks are then used as supervision ($M_{nz}\rightarrow\varphi_{nz}$) and recalibration ($M\rightarrow L_{match}$) of each other’s classifiers. This self-supervised technique encourages the main classifier $\varphi_{fg}$ to output low-value logits for noise slots and increases the magnitude difference between noise slots and foreground slots, which represent true class semantics. Meanwhile, the invalid slot mask $M_{inv}$ is constructed heuristicly (with the threshold $\alpha$) and stabilizes the early training of slot classification.
Framework Overview. As illustrated in Fig. 2, we firstly pre-train Dinosaur [36] on $\mathcal{D}_{train}$ (the training set of known classes) to obtain the representative slot features. Next, we freeze the Dinosaur part and add two classifiers, $\varphi_{nz}$ and $\varphi_{fg}$, following the encoder. After excluding the noise slots from the semantic matching, we implement the image classification. The overall loss is expressed as follows:
$L=L_{match}+L_{nz}.$ (6)
@@
Inference. For the closed-set classification, we use the max-logit slot in single-label experiments and the maximum class-wise logit across slots in multi-label experiments to determine the semantic classification result.
For OSR tasks, we score slot predictions using standard out-of-distribution (OOD) detection functions [29, 25, 15, 26]. These scores measure the degree of label shift for each slot’s associated semantic content. The final OSR decision score is obtained by summing the individual scores.
@@ V-B Results of OSR
Single-label. Tab. I compares OpenSlot with existing OSR methods, including CSSR [18], ARPL [3], MLS [41], and a transformer baseline VIT-MLS [41] First, all baselines experience a substantial accuracy degradation on the mixed test set $\mathcal{M}$ (compared to $\mathcal{H}$). This indicates that the super-label shift in the mixed OSR is more challenging than the conventional full-label shift, revealing the existing models’ limitations in processing mixed semantics.
@@ V-C Ablation studies
Effects of ANS. By comparison (Fig. 4, ours), ANS ensures that the max-logit slot falls into the FG group, aligning with the true class semantics. Moreover, ANS increases the logit value gap between the FG and Noise slot groups.
Meanwhile, we perform a similar sorting and visualization analysis of the attention masks in Fig. 6. On the right side, the main classifier $\varphi_{fg}$ suppresses noise and assigns the slot corresponding to the true class object with the largest logit value. This finding aligns with the results shown in Fig. 4 (ours). On the left, the slots sorted by the outputs of the noise classifier $\varphi_{nz}$ clearly demonstrate its ability to effectively discriminate noise features.
Fig. 4 and the comparison between Fig. 5 and 6 showcase that our ANS method effectively separates the noise slots and excludes them from semantic classification. During inference, when computing OSR scores for positive and negative images, the noise slots from both images produce similar values. This ensures that the primary score gap between positive and negative images arises from the semantic differences in class objects, facilitating a clearer class decision boundary.
Finally, we compare Pure Slot and OpenSlot in Tab. IV. Our consistent performance gains across both the $\mathcal{M}$ and $\mathcal{H}$ test sets verify that ANS enhances the ability of slots to detect both super-label and full-label shifts.
These qualitative and quantitative results demonstrate that our ANS effectively discriminates noise slot features through a two-classifier design, isolating them during classification training. This ensures that true class slots are accurately matched with their labels, mitigating semantic misalignments. As a result, OpenSlot more effectively classifies both known and unknown objects in multi-label and mixed scenarios, improving the detection of super-label shifts.
Choice of scoring functions. To obtain the OSR results, we score the slot-based class predictions. We evaluate various out-of-distribution (OOD) detection methods, including four logit-based: MSP [15], ODIN [26], Energy [29], MaxLogit [41], as well as the feature-based Mahalanobis distance [25]. In Tab. V, we compare their performances. When combined with OpenSlot, MaxLogit, and Energy [29] consistently outperform other scoring methods, yielding superior OSR results.
@@
¿m0.45cm ¿m0.45cm —m 1.4cm¡m 1.4cm¡m 1.4cm¡m 1.4cm¡ $\alpha$ $\beta$ VOC-6/14 COCO-20/60
@@ V-D Effect of slot prediction in OSR scoring
By contrast, our slot approach addresses the threat of known classes in the mixed OSR problem by producing independent class predictions for captured image semantics. During inference, every slot prediction in $f_{slot}$ is measured with the score function and is counted for OSR evaluation, without effects from others. In Fig. 7 (bottom right), we observe that the score gap between the positive and the negative image primarily comes from the detected unknown class object.
@@
V-E Yielding Benefits to Open-Set Object Detection
Results. Table VIII compares OpenSlot with existing OSOD studies. Without employing a Region Proposal Network [34], OpenSlot even outperforms several object detector-based approaches and achieves performance comparable to SIREN-vMF [8] (Detailed localization results are reported in Appendix). Fig. 8 presents example results where OpenSlot corrects the mispredictions of the detector.
Notably, unlike traditional methods that rely on predefined classes and selective search algorithms for proposal generation, OpenSlot discovers objects from image data in an unsupervised manner [2, 36], and can obtain robust object representations against domain shifts [6]. When tested on open-set images with significant domain differences from the closet-set (Fig. 9), these properties ensure that OpenSlot identifies different objects of interest, leading to improved localization of unknown objects and reflecting our slot approach’s stronger generalization ability than detector-based methods.
Overhead comparison. Tab. IX compares the computational cost of OpenSlot, Faster R-CNN (ResNet50)[34], and DETR (ResNet50)[1] in terms of model size, training time, and inference speed. Our OpenSlot features a lightweight design with only 2.4 million parameters (including the pre-training and classification stages), requiring just 3.2 hours for training. This is significantly lower than existing CNN/transformer-based object detection models. Additionally, OpenSlot demonstrates faster inference times for object detection, underscoring its efficiency advantages.
Method Param. Train. time Infer. speed
Faster-RCNN [34] 33.0 M 3.7 h 0.24 s
DETR [1] 41.6 M 3.9 h 0.29 s
Ours 2.4 M 3.2 h 0.16 s
TABLE IX: We compare Faster-RCNN and OpenSlot on RTX 4090, in terms of the parameter size (million, M), training time (hours, h), and inference speed (second/per image).